rayneeee:

4(2x – 5) = 4 How many solutions does this equation have?

CripQUEZZ:

One solution Distribute the (4) into the parentheses (4 x 2x) - (4 x -5) = 4 8x - 20 = 4 Then do the opposite of (-20) and add it to both sides 8x - 20 + 20 = 4 + 20 The (-20 + 20) cancels out and your left with 8x = 24 Then do the opposite of (8x) and divide it on both sides 8x/8 = 24/8 Then simplify and you get X = 3

rayneeee:

alr thx

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rayneeee alr thx \(\color{#0cbb34}{\text{End of Quote}}\) Np